OpenStudy (anonymous):

Can someone help me understand the two solutions to this homogenous equation?

OpenStudy (anonymous):

\[\alpha ^{n}-5\alpha ^{n-1}+6\alpha^{n-2}=0\]

OpenStudy (anonymous):

The answer is 2 and 3 how come?!